# Arithmetic Properties of Several Generalized-Constant Sequences, with Implications for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$

### v4

@@ 5. Conclusion
In the present article, we defined sequences of generalized Euler-Mascheroni, Euler-Gompertz, and Eta constants, denoted by $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$, respectively. After characterizing the basic asymptotic behavior of these sequences, we provided the following results:
$\bullet\>$ at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and – conditional on a realistic conjecture computationally verified for $2\leq n\leq 26$ – at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often, with analogous results holding for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$;
@@
Although our primary focus has been to investigate properties of the sequences $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, $\eta^{\left(n\right)}$, $\widetilde{\delta}^{\left(n\right)}$, and $\widetilde{\eta}^{\left(n\right)}$, it is important to note that further insights may be obtained by decomposing the sequence of $E$-functions in (22) by the method of multisections. For example, $F_{n}\left(z\right)$ can be split into the mod-2 multisection components
@@
each of which is itself an $E$-function, and these new functions arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$. By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $n\in\mathbb{Z}_{\geq 1}$, the $G_{0}\left(z\right),H_{0}\left(z\right),G_{1}\left(z\right),H_{1}\left(z\right),\ldots,G_{n}\left(z\right),H_{n}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z\right)$. This allows us to conclude, via the refined Siegel-Shidlovskii theorem (see Beukers [15]), that for $z=1$ the set of values $\left\{G_{0}\left(1\right),H_{0}\left(1\right),G_{1}\left(1\right),H_{1}\left(1\right),\ldots,G_{n}\left(1\right),H_{n}\left(1\right)\right\}$ are algebraically independent over $\overline{\mathbb{Q}}$, and therefore transcendental. Finally, recognizing that
@@
$e\left[-2n!\left(1+\dfrac{1}{2\cdot 3^{n+1}}+O\left(\dfrac{1}{5^{n}}\right)\right)\right]$

### v5

@@ 5. Conclusion
In the present article, we defined sequences of generalized Euler-Mascheroni, Euler-Gompertz, and Eta constants, denoted by $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$, respectively. After characterizing the basic asymptotic behavior of these sequences, we provided the following (unconditional) results:
$\bullet\>$ at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental, with analogous implications for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$;
@@
Although our primary focus has been to investigate properties of the sequences $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, $\eta^{\left(n\right)}$, $\widetilde{\delta}^{\left(n\right)}$, and $\widetilde{\eta}^{\left(n\right)}$, it is important to note that further insights may be obtained by decomposing the sequence of $E$-functions in (22) by the method of multisections. For example, $F_{n}\left(z\right),\>n\in\mathbb{Z}_{\geq 0}$ can be split into the mod-2 multisection components
@@
each of which is itself an $E$-function, and these new functions arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$. By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $m\in\mathbb{Z}_{\geq 1}$, the functions $G_{1}\left(z\right),H_{1}\left(z\right),G_{2}\left(z\right),H_{2}\left(z\right),\ldots,G_{m}\left(z\right),H_{m}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z,G_{0}\left(z\right),H_{0}\left(z\right)\right)$. This allows us to conclude, via the refined Siegel-Shidlovskii theorem (see Beukers [15]), that for $z=1$ the set of values $\left\{G_{1}\left(1\right),H_{1}\left(1\right),G_{2}\left(1\right),H_{2}\left(1\right),\ldots,G_{m}\left(1\right),H_{m}\left(1\right)\right\}$ are algebraically independent over $\overline{\mathbb{Q}}\left(G_{0}\left(1\right),H_{0}\left(1\right)\right)$, and therefore transcendental. Finally, recognizing that
@@
$n!\left(1-\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right)+\left(n!\right)o\left(\dfrac{1}{3^{n}}\right)$
@@
$e\left[-2n!\left(1+\dfrac{1}{2\cdot 3^{n+1}}+O\left(\dfrac{1}{5^{n}}\right)\right)+\left(n!\right)o\left(\dfrac{1}{5^{n}}\right)\right]$